### v1

@@ 1 Introduction
The emergence of transformer-based Large Language Models (LLMs) such as ChatGPT [27], GPT-4 [1, 18], DeepSeek [17, 50, 49, 33], and LLaMA [71, 72, 33] has driven transformative advancements across multiple domains [22]. Tasks like code generalization [47, 35], conversational systems [56, 78, 84], and mathematical reasoning [37, 82, 83], once considered exclusive to human expertise, are now routinely mastered by these AI systems.
@@
  • We first simplify the complicated matrix computation to the parallel vector computation utilizing the decoder-only property of a generative transformer-based language model. Therefore, we formalize its explicit learning dynamics, which we assume the learning behavior of LLMs is constrained to the kernel regimes due to the over-parameterization. This is referred to the Lazy Learning [42, 31], where the model merely memorizes all data points during optimization. (See Section 4 and Section 6)
@@
  • The primary contribution of this work is a two-stage upper bound on the in-distribution expected risk, which formalizes a phase transition between a compute-starved stage and a data-limited stage. Our analysis reveals that the generalization error initially undergoes an exponential decay governed by total computational expenditure, $\sf C$}, during which the expected risk is dominated by optimization dynamics. Upon reaching a critical saturation point, defined by the condition in Eq. (3), the scaling trajectory undergoes a structural shift: the rapid exponential decrease transitions into a power-law regime. In this subsequent stage, the frontier error is fundamentally constrained by the target function’s intrinsic complexity and the dataset size, $\sf N$}. Consequently, further increases in training duration or model capacity yield diminishing returns on generalization performance unless accompanied by a proportional expansion of high-quality data. (See Section 7)
@@ 2 Related Work
Scaling Law. Several recent works have empirically explored scaling laws in deep neural networks [44, 38, 60, 41, 61]. The study of neural scaling laws can also be traced back to earlier foundational works [25, 64, 8]. From a theoretical perspective, various solvable models have been developed using random feature models [20, 2, 15, 16, 59] to analyze neural scaling laws under specific limits. Additionally, theoretical analyses on linear models [75, 19, 65, 23, 54, 55] have significantly advanced our understanding of scaling laws. In contrast to these studies, our work focuses on the sequence-to-sequence stochastic training of multi-layer transformer-based language models, a topic that has not been widely discussed in prior research.
@@
Science of Transformer-based Language Models. The complex architecture and stochastic optimization processes of transformer-based language models pose significant challenges for theoretical analysis. However, developing theoretical guarantees for LLMs is essential for advancing their design and performance. Recent research has addressed various aspects of LLMs, including improving efficiency [5, 6, 7, 40, 45, 4, 29, 66], optimizing training processes [28, 52], analyzing ”white-box” transformers [80, 81, 32, 58], and investigating the emergent abilities of LLMs [22, 77, 10, 12, 11, 13]. By bridging theoretical understanding with practical advancements, these studies provide valuable insights for the development of the next generation of AI systems.
@@ 3 Preliminary
This section provides the preliminary for our analysis, where we introduce some basic notations in Section 3.1, and present the problem setup in Section 3.2. We encourage the reader to refer to Appendix A for the formal technical preliminary.
@@ 3.1 Basic Notations
Let $d]=\{1,2,\dots,d\}$. For $u\in\mathbb{R}^{d}$, the $\ell_{p}$-norm is $\|u\|_{p}:=(\sum_{k=1}^{d}|u_{k}|^{p})^{1/p}$. The Frobenius norm of $U\in\mathbb{R}^{d_{1}\times d_{2}}$ is $\|U\|_{F}:=(\sum_{(k_{1},k_{2})\in[d_{1}]\times[d_{2}]}U_{k_{1},k_{2}}^{2})^{1/2}$. For a matrix $A\in\mathbb{R}^{d\times d}$, $\lambda_{\min}(A)$ denotes its smallest eigenvalue. The indicator function $\operatorname{\mathbb{I}}\{E_{1},\dots,E_{n}\}$ equals $1$ if all events $E_{1},\dots,E_{n}$ occur, and 0 otherwise. The mapping $\operatorname{mat}:\mathbb{R}^{d^{2}}\to\mathbb{R}^{d\times d}$ reshapes a vector $a$ into a matrix such that $\operatorname{mat}_{k_{1},k_{2}}(a)=a_{(k_{1}-1)d+k_{2}}$ for $(k_{1},k_{2})\in[d]\times[d]$. Conversely, $\operatorname{vec}:\mathbb{R}^{n\times d}\to\mathbb{R}^{nd}$ flattens a matrix $A$ into a vector with $\operatorname{vec}_{k}(A)=A_{\lfloor k/d\rfloor,\,k-\lfloor k/d\rfloor\cdot d}$ for $k\in[nd]$. For a function $f:X\to\mathbb{R}^{d_{1}\times d_{2}}$, $f_{k_{1},k_{2}}(x)$ denotes the $(k_{1},k_{2})$-entry of $f(x)$. $e_{k}\in\mathbb{R}^{d}$ is the standard basis vector with a $1$ in the $k$-th entry and 0 elsewhere. Finally, we denote $a\wedge b:=\max\{a,b\}$ for $a,b\in\mathbb{R}$.
@@ 4.1 Simplifications
Utilizing the attention network’s decoder-only property, it is obvious that the matrix computation can be parallelized to vector computation. We use $X_{i,\leq\ell}\in\mathbb{R}^{\ell\times d}$ to denote the first-$\ell$ tokens ($\forall\ell\in[L]$) of matrix $X_{i}$ in $\mathbb{D}$. Hence, we compact the outputs of the model function and targets on the whole dataset to two matrices ${\sf F}(t),{\sf Y}\in\mathbb{R}^{nL\times d}$, where ${\sf F}_{(i-1)L+\ell}(t)=F_{\ell}(X_{i},\theta(t))\in\mathbb{R}^{d}$ and ${\sf Y}_{(i-1)L+\ell}=Y_{i,\ell}$ for each $(X_{i},Y_{i})$ in training dataset $\mathbb{D}$ and $\ell\in[L]$. Therefore, we derive ${\cal L}(t,\mathbb{D})=\frac{1}{n}\|{\sf F}(t)-{\sf Y}\|_{F}^{2}$ (See Lemma B.3). Besides, we list following helpful functions ($(i,\ell)\in[n]\times[L]$):
@@
  • ($\ell$-th Token of Hidden State)
    $\mu_{(\nu),(i-1)L+\ell}(t):=\frac{\omega}{\sqrt{m}}\sum_{r=1}^{m}a_{(\nu),r}\cdot\phi(\langle o_{(\nu),(i-1)L+\ell}(t),w_{(\nu),r}(t)\rangle)\in\mathbb{R}^{d},$
    where $\phi(x):=\max\{0,x\},\forall x\in\mathbb{R}$. $\mu_{(0),(i-1)L+\ell}(t)=X_{i,\ell}+E_{\ell}$.
@@ 4.2 Key Derivation for Learning Dynamics
Lemma 4.1 (Learning dynamics, informal version of Lemma B.6).
@@
Although the derivation of the gradient is complicated, the technique used for the proof is just the chain rule. We provide the complete proof in Appendix B. ∎
@@
5 Convergence of Transformers Training
In this section, we showcase the training convergence with an arbitrary error by limiting the kernel matrix $H_{(\nu)}^{\prime}(0)$ at initialization to the Neural Tangent Kernel (NTK) [42, 31] assumption. We state the formal assumption and the basic inductive proof in Section 5.1, and Section 5.2 demonstrates the results of the training convergence.
@@ 5.1 Assumptions and Inductions
Lemma 5.3 (Informal version of Lemma D.1).
@@
The stability of the PD property is ensured by $\lambda_{\min}(H^{\prime})\geq\lambda_{\min}(H)-\|H-H^{\prime}\|_{F}$ (Fact A.11) since the considerably small perturbation $\|H-H^{\prime}\|_{F}$. The upper bound on perturbation converges due to increasing feed-forward layer width $m$ and decreasing weight perturbation radius $R$, leading to the Lazy Learning regime. The complete proof of this lemma is stated in Appendix D. ∎
@@ 5.2 Convergence under Kernel Regimes
Factors of Scaling Law. To begin with, we list several crucial factors of the scaling law below. We ignore the batch size as a constant in the analysis of the total compute. In addition, the model size refers to the number of trainable parameters in our model, which could be trivially calculated, and the definition of the total compute holds as our compute analysis in Lemma E.1.
@@
Theorem 5.5 (Informal version of Theorem E.2).
@@
Firstly, we confirm the connection between hidden-state gradient norm and the training objective in Part 15 of Lemma C.1, the model convergence therefore exponentially benefits from the model size (neural depth and width) and the training time. Besides, the variance produced by the stochastic algorithm is provably reduced with a considerably large $m$. See Appendix E.2 for the detailed proofs. ∎
6 Generalization Bound of Empirical Risk Minimizer
This section formally analyzes the generalization error bound of transformer-based language models. In particular, Section 6.1 provides the approximation error w.r.t the model size $\sf M$} via the approximation capability of kernel regression in RKHS (Reproducing Kernel Hilbert Space), and Section 6.2 showcases the general scaling law, which describes how the generalization error bound converges with the growing total computational cost.
6.1 Approximation
@@
Corollary 6.1 (Informal version of Corollary E.3).
@@
According to the conditions of the target function $F^{*}$ in Section 3, we formulate its expression in the RKHS (Reproducing Kernel Hilbert Space), hence, the approximation error will be bounded by a considerably small $\varepsilon$ and over-parameterized model width $m$, which is similar to Theorem 3.2 in [3]. The formal proof is in Appendix E.3. ∎
6.2 Generalization
@@
Theorem 6.2 (Informal version of Theorem F.6).
@@
This proof utilizes the results in Lemma 11 of [63] for the generalization error and Corollary 6.1 for the approximation error. The full proof is shown in Appendix F.1. ∎
@@ 7 Scaling Law and Its Potential Drag
In this section, we unify our results from Section 5 (training convergence dynamics) and Section 6 (generalization bound of ERM) into one. We derive our theory of transformer scaling law with the discovery of a pattern transition inside in Section 7.1. Next, in Section 7.2, we discuss several vital factors that can considerably affect the success of the scaling process. For every situation that might lead to the failure of scaling, we call it potential drag of the scaling law.
@@ 7.1 General Scaling Law with Pattern Transition
Theorem 7.1 (Informal version of Theorem F.7).
@@
    $C\exp(-\alpha\xi^{2}{\sf C}/{\sf N}^{6})\cdot{\cal L}(0,\mathbb{D}),$
@@
We decompose the excess risk $\Delta{\cal R}(F)$ into three terms: i) optimization risk ${\cal R}_{\rm optimization}:=\|F(\cdot,\theta(t))-F(\cdot,\theta(\infty))\|_{L^{F}({\cal X})}^{2}$; ii) approximation risk ${\cal R}_{\rm approximation}:=\left\|F(\cdot,\theta(\infty))-{\rm arg}\inf_{F^{\prime}\in{\cal F}_{\sf M,\infty,N}}\|F^{\prime}-F^{*}\|_{L^{F}({\cal X})}^{2}\right\|_{L^{F}({\cal X})}^{2}$; and iii) estimation risk ${\cal R}_{\rm estimation}$ $\inf_{F^{\prime}\in{\cal F}_{\sf M,\infty,N}}$ $\|F^{\prime}-F^{*}\|_{L^{F}({\cal X})}^{2}$. The second and the third terms are already bounded in Theorem 6.2, and the first term relates to the training loss; we thereby utilize Theorem 5.5 to finish the proof. The formal analysis is stated in Appendix F.2. ∎
Discussion. The transition defined by the condition ${\sf C}\propto{\sf N}^{6}$ identifies a fundamental phase change in the learning trajectory, shifting the regime from one dominated by optimization dynamics to one dominated by statistical scaling. In Stage I (Compute-Starved Stage), the system is compute-starved but capacity-rich. The exponential decay with respect to $\sf C$} suggests that the model is rapidly navigating the loss landscape to find a global minimum, and in this regime, the primary constraint is the duration of training rather than the quality of the data. Conversely, Stage II (Data-Limited Stage) represents the saturation of the optimization process. Once the threshold in Eq. (3) is crossed, the model has effectively ”exhausted” the easily accessible information in the dataset. The shift to a power-law decay of ${\sf C}^{-1/6}$ signifies that further progress requires navigating the high-dimensional geometry of the target function class. This bifurcation explains why empirical scaling laws often only become observable after an initial period of training: Stage I is the transient ”discovery” phase, while Stage II represents the steady-state statistical refinement that defines modern large-scale modeling.
@@
Theorem 7.2 (Single law, informal version of Theorem F.8).
@@
This theorem is trivially derived from the intermediate results of Theorem 7.1. For three variables: $\sf T$, $\sf M$, and $\sf N$, we fix two of them to get the single law of the remaining one. The complete derivation is in Appendix F.2. ∎

### v2

@@ 1 Introduction
The emergence of transformer-based Large Language Models (LLMs) such as ChatGPT [28], GPT-4 & 5 [1, 25], DeepSeek [18, 65, 64, 41], and LLaMA [86, 87, 41] has driven transformative advancements across multiple domains [24]. Tasks like code generalization [60, 42], conversational systems [68, 94, 102], and mathematical reasoning [46, 97, 96], once considered exclusive to human expertise, are now routinely mastered by these AI systems.
@@
Figure 1: Visualization of independent variables scaling law (a.k.a single law) utilizing GPT-2 model series [24] with TinyStories dataset [38], including Model Law, Data Law, and Time Law with 2M dataset size. The corresponding theoretical analysis is in Theorem 6.1.
@@
  • We first simplify the complicated matrix computation to the parallel vector computation utilizing the decoder-only property of a generative transformer-based language model. Therefore, we formalize its explicit learning dynamics, which we assume the learning behavior of LLMs is constrained to the kernel regimes due to the over-parameterization. This is referred to as the Lazy Learning [52, 37], where the model merely memorizes all data points during optimization. (See Section 4)
@@
  • The primary contribution of this work is a two-stage upper bound on the in-distribution expected risk, which formalizes a phase transition between a compute-starved stage and a data-limited stage. Our analysis reveals that the generalization error initially undergoes an exponential decay governed by total computational expenditure, $\sf C$}, during which the expected risk is dominated by optimization dynamics. Upon reaching a critical saturation point, defined by the condition in Eq. (3), the scaling trajectory undergoes a structural shift: the rapid exponential decrease transitions into a power-law regime. In this subsequent stage, the frontier error is fundamentally constrained by the target function’s intrinsic complexity and the dataset size, $\sf N$}. Consequently, further increases in training duration or model capacity yield diminishing returns on generalization performance unless accompanied by a proportional expansion of high-quality data. We empirically substantiate these theoretical findings through a comprehensive case study on the GPT-2 series [16], which confirms the predicted phase transition between scaling regimes and validates the critical conditions under which data scaling laws break down due to noise. (See Section 6 and Section 7)
@@ 2 Related Work
Scaling Law. Several recent works have empirically explored scaling laws in deep neural networks [55, 48, 76, 47, 77, 71, 51, 74, 70, 19, 26, 44]. The study of neural scaling laws can also be traced back to earlier foundational works [27, 84, 3]. From a theoretical perspective, various solvable models have been developed using random feature models [17, 14, 15, 21, 73] to analyze neural scaling laws under specific limits. Additionally, theoretical analyses on linear models [91, 22, 79, 23, 63, 67] have significantly advanced our understanding of scaling laws. In contrast to these studies, our work focuses on the sequence-to-sequence stochastic training of multi-layer transformer-based language models, a topic that has not been widely discussed in prior research.
@@
Science of Transformer-based Language Models. The complex architecture and stochastic optimization processes of transformer-based language models pose significant challenges for theoretical analysis. However, developing theoretical guarantees for LLMs is essential for advancing their design and performance. Recent research has addressed various aspects of LLMs, including improving efficiency [10, 12, 11, 43, 54, 2, 36, 80, 62], optimizing training processes [35, 31, 59], analyzing ”white-box” transformers [98, 99, 39, 72], and investigating the emergent abilities of LLMs [24, 93, 6, 8, 7, 9]. [50] provides another theoretical treatment of generalization in transformers, focused on length extrapolation. By bridging theoretical understanding with practical advancements, these studies provide valuable insights for the development of the next generation of AI systems.
@@ 3 Preliminary
This section provides the preliminary for our analysis, where we introduce some basic notations in Section 3.1, and present the problem setup in Section 3.2. We encourage the reader to refer to Appendix B for the formal technical preliminary.
@@ 3.1 Basic Notations
Let $d]=\{1,2,\dots,d\}$. For $u\in\mathbb{R}^{d}$, the $\ell_{p}$-norm is $\|u\|_{p}:=(\sum_{k=1}^{d}|u_{k}|^{p})^{1/p}$. The Frobenius norm of $U\in\mathbb{R}^{d_{1}\times d_{2}}$ is $\|U\|_{F}:=(\sum_{(k_{1},k_{2})\in[d_{1}]\times[d_{2}]}U_{k_{1},k_{2}}^{2})^{1/2}$. For a matrix $A\in\mathbb{R}^{d\times d}$, $\lambda_{\min}(A)$ denotes its smallest eigenvalue. The indicator function $\operatorname{\mathbb{I}}\{E_{1},\dots,E_{n}\}$ equals $1$ if all events $E_{1},\dots,E_{n}$ occur, and 0 otherwise. The mapping $\operatorname{mat}:\mathbb{R}^{d^{2}}\to\mathbb{R}^{d\times d}$ reshapes a vector $a$ into a matrix such that $\operatorname{mat}_{k_{1},k_{2}}(a)=a_{(k_{1}-1)d+k_{2}}$ for $(k_{1},k_{2})\in[d]\times[d]$. Conversely, $\operatorname{vec}:\mathbb{R}^{n\times d}\to\mathbb{R}^{nd}$ flattens a matrix $A$ into a vector with $\operatorname{vec}_{k}(A)=A_{\lfloor k/d\rfloor,\,k-\lfloor k/d\rfloor\cdot d}$ for $k\in[nd]$. For a function $f:X\to\mathbb{R}^{d_{1}\times d_{2}}$, $f_{k_{1},k_{2}}(x)$ denotes the $(k_{1},k_{2})$-entry of $f(x)$. $e_{k}\in\mathbb{R}^{d}$ is the standard basis vector with a $1$ in the $k$-th entry and 0 elsewhere.
@@ 4.1 Simplifications
Utilizing the attention network’s decoder-only property, it is obvious that the matrix computation can be parallelized to vector computation. We use $X_{i,\leq\ell}\in\mathbb{R}^{\ell\times d}$ to denote the first-$\ell$ tokens ($\forall\ell\in[L]$) of matrix $X_{i}$ in $\mathbb{D}$. Hence, we compact the outputs of the model function and targets on the whole dataset to two matrices ${\sf F}(t),{\sf Y}\in\mathbb{R}^{nL\times d}$, where ${\sf F}_{(i-1)L+\ell}(t)=F_{\ell}(X_{i},\theta(t))\in\mathbb{R}^{d}$ and ${\sf Y}_{(i-1)L+\ell}=Y_{i,\ell}$ for each $(X_{i},Y_{i})$ in training dataset $\mathbb{D}$ and $\ell\in[L]$.
Therefore, we derive ${\cal L}(t,\mathbb{D})=\frac{1}{n}\|{\sf F}(t)-{\sf Y}\|_{F}^{2}$ (See Lemma C.3). Besides, we list following helpful functions ($(i,\ell)\in[n]\times[L]$):
@@ 4.2 Key Derivation for Learning Dynamics
Lemma 4.1 (Learning dynamics, informal version of Lemma C.6).
@@
Although the derivation of the gradient is complicated, the technique used for the proof is just the chain rule. We provide the complete proof in Appendix C. ∎
@@
5 Training Convergence, Approximation, and Generalization
In this section, we showcase the training convergence with an arbitrary error by limiting the kernel matrix $H_{(\nu)}^{\prime}(0)$ at initialization to the Neural Tangent Kernel (NTK) [52, 37] assumption. We state the formal assumption and the basic inductive proof in Section 5.1, and Section 5.2 demonstrates the results of the training convergence. Finally, Section 5.3 provides the approximation error w.r.t the model size $\sf M$} via the approximation capability of kernel regression in RKHS (Reproducing Kernel Hilbert Space) and the generalization upper bound of the ideal empirical risk minimizer.
@@ 5.1 Assumptions and Inductions
Comments on Definition 5.2. While the scaling requirements in Definition 5.2 may appear distinct from standard heuristic initializations, they serve as essential theoretical proxies for ensuring stability in the NTK regime. Specifically, the choice of $\omega$ mirrors the zero-output or small-norm initialization assumptions common in deep learning theory or training super-deep transformers [16, 101, 90], effectively mitigating the residual branch’s impact to guarantee a minimum input norm for each layer. Furthermore, regarding the model width $m$, we highlight the inverse dependence on depth ($m\propto N^{-2}$). This relationship suggests a trade-off where deeper architectures alleviate the strict lower bound on width, implying that the total parameter capacity can be distributed toward depth to satisfy the condition ${\cal F}_{\sf M,T,N}$ without requiring an exponentially prohibitive width.
Lemma 5.3 (Informal version of Lemma E.1).
@@
The stability of the PD property is ensured by $\lambda_{\min}(H^{\prime})\geq\lambda_{\min}(H)-\|H-H^{\prime}\|_{F}$ (Fact B.11) since the considerably small perturbation $\|H-H^{\prime}\|_{F}$. The upper bound on perturbation converges due to increasing feed-forward layer width $m$ and decreasing weight perturbation radius $R$, leading to the Lazy Learning regime. The complete proof of this lemma is stated in Appendix E. ∎
@@ 5.2 Convergence under Kernel Regimes
Factors of Scaling Law. To begin with, we list several crucial factors of the scaling law below. We ignore the batch size as a constant in the analysis of the total compute. In addition, the model size refers to the number of trainable parameters in our model, which could be trivially calculated, and the definition of the total compute holds as our compute analysis in Lemma F.1.
@@
Theorem 5.5 (Informal version of Theorem F.2).
@@
Firstly, we confirm the connection between hidden-state gradient norm and the training objective in Part 15 of Lemma D.1, the model convergence therefore exponentially benefits from the model size (neural depth and width) and the training time. Besides, the variance produced by the stochastic algorithm is provably reduced with a considerably large $m$. See Appendix F.2 for the detailed proofs. ∎
5.3 Approximation and Generalization of ERM
@@
Corollary 5.6 (Informal version of Corollary F.3).
@@
According to the conditions of the target function $F^{*}$ in Section 3, we formulate its expression in the RKHS (Reproducing Kernel Hilbert Space), hence, the approximation error will be bounded by a considerably small $\varepsilon$ and over-parameterized model width $m$, which is similar to Theorem 3.2 in [13]. The formal proof is in Appendix F.3. ∎
@@
Theorem 5.7 (Informal version of Theorem G.6).
@@
This proof utilizes the results in Lemma 11 of [78] for the generalization error and Corollary 5.6 for the approximation error. The full proof is shown in Appendix G.1. ∎
@@ 6 Scaling Law and Its Potential Drag
In this section, we unify our results from Section 5 - training convergence dynamics and generalization bound of ERM - into one. We derive our theory of transformer scaling law with the discovery of a pattern transition inside in Section 6.1. Next, in Section 6.2, we discuss several vital factors that can considerably affect the success of the scaling process. For every situation that might lead to the failure of scaling, we call it potential drag of the scaling law.
@@ 6.1 General Scaling Law with Pattern Transition
Theorem 6.1 (Informal version of Theorem G.7).
@@
    $\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq C\exp(-\alpha\xi^{2}{\sf C}/{\sf N}^{6})\cdot{\cal L}(0,\mathbb{D}),$
@@
    $\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq$ $~\xi^{5/3}\cdot O\left(\frac{{\sf C}}{W({\sf C}/\xi^{10})}\right)^{-1/6}$
    $\leq$ $~\xi^{5/3}\cdot\widetilde{O}\left({\sf C}\right)^{-1/6},$
@@
We decompose the excess risk $\Delta{\cal R}(F)$ into three terms: i) optimization risk ${\cal R}_{\rm optimization}:=\|F(\cdot,\theta(t))-F(\cdot,\theta(\infty))\|_{L^{F}({\cal X})}^{2}$; ii) approximation risk ${\cal R}_{\rm approximation}:=\left\|F(\cdot,\theta(\infty))-{\rm arg}\inf_{F^{\prime}\in{\cal F}_{\sf M,\infty,N}}\|F^{\prime}-F^{*}\|_{L^{F}({\cal X})}^{2}\right\|_{L^{F}({\cal X})}^{2}$; and iii) estimation risk ${\cal R}_{\rm estimation}$ $\inf_{F^{\prime}\in{\cal F}_{\sf M,\infty,N}}$ $\|F^{\prime}-F^{*}\|_{L^{F}({\cal X})}^{2}$. The second and the third terms are already bounded in Theorem 5.7, and the first term relates to the training loss; we thereby utilize Theorem 5.5 to finish the proof. The formal analysis is stated in Appendix G.2. ∎
Discussion. The transition defined by the condition ${\sf C}\propto{\sf N}^{6}$ identifies a fundamental phase change in the learning trajectory, shifting the regime from one dominated by optimization dynamics to one dominated by statistical scaling. In Stage I (Compute-Starved Stage), the system is compute-starved but capacity-rich. The exponential decay with respect to $\sf C$} suggests that the model is rapidly navigating the loss landscape to find a global minimum, and in this regime, the primary constraint is the duration of training rather than the quality of the data. Conversely, Stage II (Data-Limited Stage) represents the saturation of the optimization process. Once the threshold in Eq. (3) is crossed, the model has effectively exhausted the easily accessible information in the dataset. The shift to a power-law decay of ${\sf C}^{-1/6}$ signifies that further progress requires navigating the high-dimensional geometry of the target function class. This bifurcation explains why empirical scaling laws often only become observable after an initial period of training: Stage I is the transient discovery phase, while Stage II represents the steady-state statistical refinement that defines modern large-scale modeling.
@@
Theorem 6.2 (Single law, informal version of Theorem G.8).
@@
This theorem is trivially derived from the intermediate results of Theorem 6.1. For three variables: $\sf T$, $\sf M$, and $\sf N$, we fix two of them to get the single law of the remaining one. The complete derivation is in Appendix G.2. ∎
Figure 2: Scaling law curves illustrating validation of Cross-Entropy Loss across varying compute scales. The left panel utilizes a linear x-axis to highlight absolute compute expenditure, while the middle and right panels employ logarithmic scales to visualize the specific decay rates of loss. The data distinguishes between different operational regimes, specifically identifying a Compute-Starved Stage and a Data-Limited Stage. To provide a granular view of performance, markers use distinct colors and shapes to differentiate between model sizes (ranging from 124M to 1.5B parameters) and dataset sizes. These visualizations specifically compare Time-Dominated exponent-like patterns against Data-Dominated linear-like patterns, demonstrating how scaling strategies affect the capture of underlying data patterns.